OpenStudy (anonymous):

find the coordinates of the missing endpoint if P is the midpoint of NQ N(5,8), P(4,3) please help!

OpenStudy (anonymous):

midpoint of a line can be found using\[x_P=\frac{x_N+x_Q}{2}\]\[y_P=\frac{y_N+y_Q}{2}\]plug-in the coordinates of given points then solve for Q(\(x_Q,y_Q\))...